# Single-loop Algorithms for Stochastic Non-convex Optimization with Weakly-Convex Constraints

### v1

@@ Appendix A More Experiment Results
The experiments of AUC Maximization with ROC Fairness Constraints in our paper is run on a GPU server with 10 A30 24G GPUs. The experiment of continual learning with non-forgetting constraints is run on High Performance Research Computing platforms, which contains 117 GPU nodes, each with two A100 40GB GPUs.
We give other experiment results in this Appendix. First, we compute the minimum singular value $\delta$ of the Jacobian matrix of violating constraint functions in our experiments at different solutions to verify the conditions in Lemma 4.4, as shown in the Table 2. For the first experiment, we compute $\sigma$ based on the Adult dataset at the initial model, final model, and two randomly selected intermediate epochs. For our second experiment, we compute $\sigma$ when Target Foggy at both the initial and final models. All results demonstrate that the minimum singular value $\sigma$ of $H(\mathbf{x})$ remains positive, which is consistent with the theoretical guarantee stated in Lemma 4.4.
@@
Minimum Singular Values $\sigma$ of $\nabla H(\mathbf{x})$ 0.000648 0.000112 0.000102 0.000101
@@
Minimum Singular Values $\delta$ of $H(\mathbf{x})$ 24.1038 16.3397
Table 2: Minimum singular values $\delta$ of the Jacobian matrix $H(\mathbf{x})$ at different solution.
@@
Figure 5: Training curves of 5 constraint values in zero-one loss of different methods for continual learning with non-forgetting constraints when targeting the overcast class. Top: hinge penalty method with different $\beta$; Bottom: squared-hinge penalty method with different $\beta$.
Figure 6: Training curves of 5 constraint values in zero-one loss of different methods for continual learning with non-forgetting constraints when targeting the tunnel class. Top: hinge penalty method with different $\beta$; Bottom: squared-hinge penalty method with different $\beta$.
@@ Appendix B Technique Lemmas
[Lemma 1 [20]] Consider update (15). Under Assumptions 3.1, 5.1 and 5.4, by setting $\gamma_{1}^{\prime}=\frac{n-|\mathcal{B}|}{|\mathcal{B}|(1-\gamma_{1})}+(1-\gamma_{1})$, for $\gamma_{1}\leq\frac{1}{2}$, the function value variance $\Xi^{t+1}:=\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{t+1}-g_{i}(\mathbf{x}_{t+1})\|^{2}$ can be bounded as
@@
[Lemma 1 [20]] Consider update (12). Under Assumptions 3.1 and 5.1, by setting $\gamma_{2}^{\prime}=\frac{m-|\mathcal{B}_{c}|}{|\mathcal{B}_{c}|(1-\gamma_{2})}+(1-\gamma_{2})$, for $\gamma_{2}\leq\frac{1}{2}$, the function value variance $\Gamma^{t+1}:=\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{t+1}-h_{k}(\mathbf{x}_{t+1})\|^{2}$ can be bounded as
@@
From the analysis of Lemma 4.5 in [16], the following two inequalities hold true based on Lemma B.1 and Lemma B.2. For simplicity, we use a constant $M^{2}\geq\max\{2L_{f}^{2}L_{g}^{2}+2\beta^{2}L_{h}^{2},2\sigma_{f}^{2}+2L_{F}^{2}+2\beta^{2}L_{h}^{2}\}$.
@@
Notice that we have assumed $F(\cdot)$ to be $\rho_{0}$-weakly convex in Assumption 3.1. By the definition of weak convexity, it suffices to show $[h_{k}(\mathbf{x})]_{+}$ to be $\rho_{1}$-weakly convex for all $k=1,\dots,m$.
Let $[h_{k}(x)]^{\prime}_{+}$ denote any element in $\partial[h_{k}(x)]_{+}$. Given any $k\in\{1,\dots,m\}$ and $x,x^{\prime}$, following from the convexity of $[\cdot]_{+}$, we have
@@
where $(a)$ follows from the fact $[h_{k}(x^{\prime})]^{\prime}_{+}\geq 0$ for all $x^{\prime}$ and the $\rho_{1}$-weak convexity of $h_{k}(\cdot)$, and $(b)$ follows from the fact $[h_{k}(x^{\prime})]^{\prime}_{+}\leq 1$.
@@
Our goal is to prove that there exists a constant $\delta>0$ for any $\mathbf{x}$ such that $h(\mathbf{x})>0$ (violating the constraint), we have $\|\nabla h(\mathbf{x})\|\geq\delta$.
@@
$c<h(\mathbf{x})-\min_{\mathbf{y}}h(\mathbf{y})\leq\frac{\|\nabla h(\mathbf{x})\|^{2}}{2\mu}.$
This means $\|\nabla h(\mathbf{x})\|\geq\sqrt{2c\mu}=:\delta$. ∎
@@
Let $\boldsymbol{\xi}=(\xi_{1},\ldots,\xi_{m})^{T}$, where $\xi_{k}$ defined as
@@
Then
$\left\{\sum_{k=1}^{m}\xi_{k}\nabla h_{k}(\mathbf{x})\right\}=\left\{\nabla H(\mathbf{x})^{\top}\boldsymbol{\xi}\right\}\subseteq\sum_{k=1}^{m}\partial[h_{k}(\mathbf{\mathbf{x}})]_{+},$
where $\nabla H(\mathbf{x})=(\nabla h_{1}(\mathbf{x}),\ldots,\nabla h_{k}(\mathbf{x}))^{T}.$ Then for $\forall\mathbf{x}\in\mathcal{V}$, where $\mathcal{V}=\{\mathbf{x}:\max_{k}h_{k}(\mathbf{x})>0\},$
$\operatorname{dist}\left(0,\frac{1}{m}\sum_{k=1}^{m}\partial[h_{k}(\mathbf{x})]_{+}\right)\geq\left\|\frac{1}{m}\nabla H(\mathbf{x})^{\top}\boldsymbol{\xi}\right\|\geq\frac{1}{m}\lambda_{\min}(\nabla H(\mathbf{x}))\|\boldsymbol{\xi}\|\geq\frac{\sigma}{m}.$
@@
where the first inequality uses the convexity of function $[\cdot]_{+}$, the second inequality uses the fact $[u_{2,k}^{t}]^{\prime}_{+}\geq 0$ and the $\rho_{1}$-weak convexity of $h_{k}(\cdot)$, and the last inequality uses the fact that $0\leq[u_{2,k}^{t}]^{\prime}_{+}\leq 1$. Then it follows
@@
Adding above two estimation (26) and (27) back to (25), we have
@@
Plugging inequality (29) to (28), we get
@@
Taking the full expectation on both sides of (30) and applying the inequality of (20), we have
@@
Taking summation from $t=0$ to $T-1$ yields
@@
Appendix D Proof of Theorem 5.5 and D.1
First, we present the Algorithm 2 in this Appendix.
Theorem D.1.
Suppose Assumption 3.1, 5.1 and 5.4 with condition(ii) hold. Let $\gamma=\mathcal{O}(\min\{\frac{|\mathcal{B}_{2,k}|}{\beta^{4}}\epsilon^{4},\frac{|\mathcal{B}_{1,i}|}{\beta}\epsilon^{2}\})$, $\eta=\mathcal{O}(\min\{\frac{|\mathcal{B}||\mathcal{B}_{1,i}|^{1/2}\epsilon^{2}}{n\beta^{2}},\frac{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{4}}{\beta^{5}m}\})$. After $T=\mathcal{O}(\max\{\frac{m\beta^{6}}{|\mathcal{B}_{2,k}|^{1/2}|\mathcal{B}_{c}|\epsilon^{6}},\frac{n\beta^{3}}{|\mathcal{B}||\mathcal{B}_{1,i}|^{1/2}\epsilon^{4}},\frac{n\beta^{2}}{|\mathcal{B}||\mathcal{B}_{1,i}|\epsilon^{4}}\})$ iterations, the Algorithm 2 satisfies $\mathbb{E}[\|\nabla\Phi_{\lambda}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\lambda=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$
@@
To this end, we first give the proof of Theorem 5.5.
Algorithm 2 Algorithm for solving (3) under Setting II
@@
2:  for $t=0$ to $T-1$ do
3:     Sample $\mathcal{B}^{t}\in\{1,\ldots,n\}$, $\mathcal{B}^{t}_{c}\in\{1,\ldots,m\}$
@@
6:        Update $u_{1,i}^{t}$ by Eqn.(15)
@@
9:     Compute $G_{1}^{t}$ in Eqn (16)
@@
12:        Update the estimator $u_{2,k}^{t+1}$ by Eqn (12)
@@
15:     Compute $G_{2}^{t}$ as in Eqn (13)
@@
Next we bound $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle$. For given $k\in\{1,\ldots,m\}$, from (27), we have
@@
Combining inequality (31), (32) and (33) yields
@@
Plugging inequalities (35) to (34), we get
@@
Taking the full expectation on both sides of (36) and adding the result of Lemma B.1, (21) and (20), we have
@@
Taking summation from $t=0$ to $T-1$ yields
@@
[…]
@@
From (31), also denote $\bar{\mathbf{x}}_{t}:=\text{prox}_{\Phi/\bar{\rho}}(\mathbf{x}_{t})$ and consider change in the Moreau envelope, we have
@@
Adding above two estimation to (37), we have
@@
Adding above inequality to (38), we have
@@
Adding the result of Lemma B.1 and inequality (20), we get
@@
Taking summation from $t=0$ to $T-1$ yields
@@
[…]

### v2

@@ Appendix A More Experiment Results
The experiments of AUC Maximization with ROC Fairness Constraints in our paper is run on a GPU server with 10 A30 24G GPUs. Each seed in this experiment takes about 4 hours to complete. The experiment of continual learning with non-forgetting constraints is run on two A100 40GB GPUs. Each seed in this setting takes approximately 12 hours to finish. These runtimes are comparable across all methods since they share identical model architectures and computational setups, including the learning-rate schedule, weight decay, batch sizes, and the selections of $|\mathcal{B}_{c}|$ and $|\mathcal{B}_{k}|$. We tune the penalty coefficient $\beta$, which allows us to study how different methods reduce constraint violations and how their performances in the objective value change when the constraints are satisfied. This consistent setup provides a fair and controlled comparison of the optimization behavior of all methods.
We give other experiment results in this Appendix. First, we compute the minimum singular value $\delta$ of the Jacobian matrix of violating constraint functions in our experiments at different solutions to verify the conditions in Lemma 4.4, as shown in the Table 2. For the first experiment, we compute $\sigma$ based on the Adult dataset at the initial model, final model, and two randomly selected intermediate epochs. For our second experiment, we compute $\sigma$ when Target Foggy at both the initial and final models. All results demonstrate that the minimum singular value $\sigma$ of $\partial H(\mathbf{x})$ remains positive, which is consistent with the theoretical guarantee stated in Lemma 4.4.
@@
Minimum Singular Values $\sigma$ of $\partial H(\mathbf{x})$ 0.000648 0.000112 0.000102 0.000101
@@
Minimum Singular Values $\sigma$ of $\partial H(\mathbf{x})$ 24.1038 16.3397
Table 2: Minimum singular values $\sigma$ of the Jacobian matrix $\partial H(\mathbf{x})$ at different solution.
@@
Figure 5: Training curves of 4 constraint values in zero-one loss of different methods for continual learning with non-forgetting constraints when targeting the overcast class. Top: hinge penalty method with different $\beta$; Bottom: squared-hinge penalty method with different $\beta$.
Figure 6: Training curves of 4 constraint values in zero-one loss of different methods for continual learning with non-forgetting constraints when targeting the tunnel class. Top: hinge penalty method with different $\beta$; Bottom: squared-hinge penalty method with different $\beta$.
@@ Appendix B Technique Lemmas
The following lemma follows from standard result (Davis and Drusvyatskiy, 2019).
@@
Lemma B.2.
[Lemma 1 (Jiang et al., 2022)] Consider update (26). Under Assumptions 3.1, 5.1 and D.1, by setting $\gamma_{1}^{\prime}=\frac{n-|\mathcal{B}|}{|\mathcal{B}|(1-\gamma_{1})}+(1-\gamma_{1})$, for $\gamma_{1}\leq\frac{1}{2}$, the function value variance $\Xi^{t+1}:=\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{t+1}-g_{i}(\mathbf{x}_{t+1})\|^{2}$ can be bounded as
@@
[Lemma 1 (Jiang et al., 2022)] Consider the update Line 6 in Algorithm 1. Under Assumptions 3.1 and 5.1, by setting $\gamma_{2}^{\prime}=\frac{m-|\mathcal{B}_{c}|}{|\mathcal{B}_{c}|(1-\gamma_{2})}+(1-\gamma_{2})$, for $\gamma_{2}\leq\frac{1}{2}$, the function value variance $\Gamma^{t+1}:=\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{t+1}-h_{k}(\mathbf{x}_{t+1})\|^{2}$ can be bounded as
@@
From the analysis of Lemma 4.5 in Hu et al. (2024), the following two inequalities hold true based on Lemma B.2 and Lemma B.3. For simplicity, we use a constant $M^{2}\geq\max\{2L_{f}^{2}L_{g}^{2}+2\beta^{2}L_{h}^{2},2\sigma_{f}^{2}+2L_{F}^{2}+2\beta^{2}L_{h}^{2}\}$.
@@
Notice that we have assumed $F(\cdot)$ to be $\rho_{0}$-weakly convex in Assumption 3.1. By the definition of weak convexity, it suffices to show $h_{k}(\mathbf{x})]_{+}$ to be $\rho_{1}$-weakly convex for all $k=1,\dots,m$.
Let $h_{k}(x)]^{\prime}_{+}$ denote any element in $\partial[h_{k}(x)]_{+}$. Given any $k\in\{1,\dots,m\}$ and $x,x^{\prime}$, following from the convexity of $\cdot$$_{+}$, we have
@@
where $(a)$ follows from the fact $h_{k}(x^{\prime})]^{\prime}_{+}\geq 0$ for all $x^{\prime}$ and the $\rho_{1}$-weak convexity of $h_{k}(\cdot)$, and $(b)$ follows from the fact $h_{k}(x^{\prime})]^{\prime}_{+}\leq 1$.
@@
Our goal is to prove that there exists a constant $\delta>0$ for any $\mathbf{x}$ such that $h(\mathbf{x})>0$ (violating the constraint), we have $\text{dist}(0,\partial h(\mathbf{x}))\geq\delta$.
@@
$c<h(\mathbf{x})-\min_{\mathbf{y}}h(\mathbf{y})\leq\frac{\text{dist}(0,\partial h(\mathbf{x}))^{2}}{2\mu}.$
This means $\|\text{dist}(0,\partial h(\mathbf{x}))|\geq\sqrt{2c\mu}=:\delta$. ∎
@@
Let $\bm{\xi}=(\xi_{1},\ldots,\xi_{m})^{T}$, where $\xi_{k}$ defined as
@@
For any $\mathbf{x}$ and any sub-Jacobian $G(\mathbf{x})\in\partial H(\mathbf{x})$, let $g_{k}(\mathbf{x})^{\top}$ denote its $k$-th row (so $g_{k}(\mathbf{x})\in\partial h_{k}(\mathbf{x})$).
$\Bigl\{\sum_{k=1}^{m}\xi_{k}g_{k}(\mathbf{x})\Bigr\}\;=\;\{G(\mathbf{x})^{\top}\bm{\xi}\}\;\subset\;\sum_{k=1}^{m}\partial[h_{k}(\mathbf{x})]_{+}.$
Let $\mathcal{V}:=\{\mathbf{x}:\max_{k}h_{k}(\mathbf{x})>0\}$. For any $\mathbf{x}\in\mathcal{V}$ we have
$\operatorname{dist}\!\left(0,\frac{1}{m}\sum_{k=1}^{m}\partial[h_{k}(\mathbf{x})]_{+}\right)\;\geq\;\left\|\frac{1}{m}G(\mathbf{x})^{\top}\bm{\xi}\right\|\;\geq\;\frac{1}{m}\,\lambda_{\min}\bigl(G(\mathbf{x})\bigr)\,\|\bm{\xi}\|\;\geq\;\frac{1}{m}\,\lambda_{\min}\bigl(G(\mathbf{x})\bigr).$
By the assumption that $\text{dist}(0,\lambda_{\min}(\partial H(\mathbf{x})))\geq\sigma>0,\forall\mathbf{x}\text{ such that }\max_{k}h_{k}(\mathbf{x})>0$, we have
$\lambda_{\min}\bigl(G(\mathbf{x})\bigr)\;\geq\;\sigma\quad\forall\,G(\mathbf{x})\in\partial H(\mathbf{x}).$
Therefore, for all $\mathbf{x}\in\mathcal{V}$,
$\operatorname{dist}\!\left(0,\frac{1}{m}\sum_{k=1}^{m}\partial[h_{k}(\mathbf{x})]_{+}\right)\;\geq\;\frac{\sigma}{m}$
@@
where the first inequality uses the convexity of function $\cdot$$_{+}$, the second inequality uses the fact $u_{2,k}^{t}]^{\prime}_{+}\geq 0$ and the $\rho_{1}$-weak convexity of $h_{k}(\cdot)$, and the last inequality uses the fact that $0\leq[u_{2,k}^{t}]^{\prime}_{+}\leq 1$. Then it follows
@@
Adding above two estimation equation 19 and equation 20 back to equation 18, we have
@@
Plugging inequality equation 22 to equation 21, we get
@@
Taking the full expectation on both sides of equation 23 and applying the inequality of equation 13, we have
@@
Taking summation from $t=0$ to T1T-1 yields
@@
Now, we will prove that $\mathbf{x}_{\hat{t}}$ is a nearly $\epsilon$-KKT solution, where (i) holds in expectation, (ii) and (iii) in Definition 3.3 hold in high probability, to the original problem (1).
@@
[…]
$0\in\partial F(\bar{\mathbf{x}}_{\hat{t}})+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}}_{\hat{t}})+(\bar{\mathbf{x}}_{\hat{t}}-{\mathbf{x}}_{\hat{t}})\bar{\rho}.$
Since $\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{\hat{t}})\|]\leq\mathcal{O}(\epsilon)$, then we have $\mathbb{E}[\|\bar{\mathbf{x}}_{\hat{t}}-{\mathbf{x}}_{\hat{t}}\|]=\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{x}_{\hat{t}})\|]/\bar{\rho}\leq\epsilon/\bar{\rho}$. Then we have
$\mathbb{E}\Big[\text{dist}\big(0,\partial F(\bar{\mathbf{x}}_{\hat{t}})+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}}_{\hat{t}})\big)\Big]\leq\mathbb{E}[\|(\bar{\mathbf{x}}_{\hat{t}}-{\mathbf{x}_{\hat{t}}})\bar{\rho}\|]\leq\epsilon.$
[…]
$\text{dist}\left(0,\mathbf{v}+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}}_{\hat{t}})\right)\geq\text{dist}\left(0,\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}}_{\hat{t}})\right)-\|\mathbf{v}\|\geq\beta\delta-L_{F}\geq 0.$ (24)
Therefore,
$\epsilon$ $)]\displaystyle\geq\mathbb{E}\Big[\text{dist}\big(0,\mathbf{v}+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}}_{\hat{t}}$$\big)\Big$]
$=\mathbb{E}\Big[\text{dist}\big(0,\mathbf{v}+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}}_{\hat{t}})\big)\big|\max_{k}h_{k}(\bar{\mathbf{x}}_{\hat{t}})>\epsilon\Big]\text{Prob}(\max_{k}h_{k}(\bar{\mathbf{x}}_{\hat{t}})>\epsilon)$
$+\mathbb{E}\Big[\text{dist}\big(0,\mathbf{v}+\frac{\beta}{m}\sum\nolimits_{k=1}^{m}\partial h_{k}^{+}(\bar{\mathbf{x}}_{\hat{t}})\big)\big|\max_{k}h_{k}(\bar{\mathbf{x}}_{\hat{t}})\leq\epsilon\Big]\text{Prob}(\max_{k}h_{k}(\bar{\mathbf{x}}_{\hat{t}})\leq\epsilon)$
$\geq\text{Prob}(\max_{k}h_{k}(\bar{\mathbf{x}}_{\hat{t}})>\epsilon)(\beta\delta-L_{F}).$
As a result,
$\text{Prob}(\max_{k}h_{k}(\bar{\mathbf{x}}_{\hat{t}})>\epsilon)\leq\frac{\epsilon}{\beta\delta-L_{F}}=\mathcal{O}(\epsilon)$ (25)
[…]
Appendix D Analysis of Setting II
Assumption D.1.
[…]
Then we update $\mathbf{x}_{t+1}$ similarly as before by $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta(G_{1}^{t}+G_{2}^{t})$. Due to the limit of space, we present full steps in Algorithm 2.
Algorithm 2 Algorithm for solving equation 2 under Setting II
@@ Appendix D Analysis of Setting II
2:for $t=0$ to T1T-1 do
3:  Sample $\mathcal{B}^{t}\subset\{1,\ldots,n\}$, $\mathcal{B}^{t}_{c}\subset\{1,\ldots,m\}$
@@
6:   Update $u_{1,i}^{t}$ by $u_{1,i}^{t+1}=(1-\gamma_{1})u_{1,i}^{t}+\gamma_{1}{g_{i}}(\mathbf{x}_{t};\mathcal{B}_{1,i}^{t})+\gamma_{1}^{\prime}({g_{i}}(\mathbf{x}_{t};\mathcal{B}_{1,i}^{t})-{g_{i}}(\mathbf{x}_{t-1};\mathcal{B}_{1,i}^{t})),$
@@
9:  Compute $G_{1}^{t}=$ $\frac{1}{|\mathcal{B}^{t}|}\sum_{i\in\mathcal{B}^{t}}\partial g_{i}(\mathbf{x}_{t};\mathcal{B}_{1,i}^{t})\partial f_{i}(u_{1,i}^{t}).$
@@
12:   Update $u_{2,k}^{t+1}=(1-\gamma_{2})u_{2,k}^{t}+\gamma_{2}{h_{k}}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})+\gamma_{2}^{\prime}({h_{k}}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})-{h_{k}}(\mathbf{x}_{t-1};\mathcal{B}_{2,k}^{t}))$
@@
15:  Compute $G_{2}^{t}=\frac{\beta}{|\mathcal{B}^{t}_{c}|}\sum_{k\in\mathcal{B}^{t}_{c}}\partial h_{k}(\mathbf{x}_{t};\mathcal{B}_{2,k}^{t})[u_{2,k}^{t}]^{\prime}_{+}$, $G_{1}^{t}=\frac{1}{|\mathcal{B}^{t}|}\sum_{\zeta\in\mathcal{B}^{t}}\partial f(\mathbf{x}_{t};\zeta)$.
@@
The convergence of Algorithm 2 under condition (i) of Assumption D.1 is stated in Theorem D.2, and that under condition (ii) of Assumption D.1 is stated in Theorem D.3.
Theorem D.2.
Suppose Assumption 3.1, 5.1, D.1 with condition (i) hold. Let $\gamma_{1}=\gamma_{2}=\mathcal{O}(\min\{|\mathcal{B}_{1,i}|,\frac{|\mathcal{B}_{2,k}|}{\beta^{2}}\}\frac{\epsilon^{4}}{\beta^{2}})$, $\eta=\mathcal{O}(\min\{\frac{|\mathcal{B}|}{n},\frac{|\mathcal{B}_{c}|}{\beta m}\}\min\{|\mathcal{B}_{1,i}|^{1/2},\frac{|\mathcal{B}_{2,k}|^{1/2}}{\beta}\}\frac{\epsilon^{4}}{\beta^{3}})$. After $T=\mathcal{O}(\max\{\frac{\beta}{|\mathcal{B}_{1,i}|^{1/2}},\frac{\beta^{2}}{|\mathcal{B}_{2,k}|^{1/2}},\frac{1}{|\mathcal{B}_{1,i}|}\}\max\{\frac{n}{|\mathcal{B}|},\frac{\beta m}{|\mathcal{B}_{c}|}\}\frac{\beta^{3}}{\epsilon^{6}})$ iterations, the Algorithm 2 satisfies $\mathbb{E}[\|\nabla\Phi_{\theta}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\theta=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$
Remark: Combining the above result and that in Theorem 4.2, Algorithm 2 needs $T=\mathcal{O}(\frac{1}{\epsilon^{6}})$ to achieve a nearly $\epsilon$-KKT solution to the original problem equation 1. It is better than the complexity of $\mathcal{O}(1/\epsilon^{7})$ of the single-loop algorithm considered in Li et al. (2024a).
Theorem D.3.
Suppose Assumption 3.1, 5.1 and D.1 with condition (ii) hold. Let $\gamma=\mathcal{O}(\min\{\frac{|\mathcal{B}_{2,k}|}{\beta^{4}}\epsilon^{4},\frac{|\mathcal{B}_{1,i}|}{\beta}\epsilon^{2}\})$, $\eta=\mathcal{O}(\min\{\frac{|\mathcal{B}||\mathcal{B}_{1,i}|^{1/2}\epsilon^{2}}{n\beta^{2}},\frac{|\mathcal{B}_{c}||\mathcal{B}_{2,k}|^{1/2}\epsilon^{4}}{\beta^{5}m}\})$. After $T=\mathcal{O}(\max\{\frac{m\beta^{6}}{|\mathcal{B}_{2,k}|^{1/2}|\mathcal{B}_{c}|\epsilon^{6}},\frac{n\beta^{3}}{|\mathcal{B}||\mathcal{B}_{1,i}|^{1/2}\epsilon^{4}},\frac{n\beta^{2}}{|\mathcal{B}||\mathcal{B}_{1,i}|\epsilon^{4}}\})$ iterations, the Algorithm 2 satisfies $\mathbb{E}[\|\nabla\Phi_{\lambda}(\mathbf{x}_{\hat{t}})\|]\leq\epsilon$ from some $\lambda=O(1/(\rho_{0}+\rho_{1}\beta))$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$
Proof of Theorem D.2.
@@
Next we bound $\langle\bar{\mathbf{x}}_{t}-\mathbf{x}_{t},\mathbb{E}_{t}[G_{2}^{t}]\rangle$. For given $k\in\{1,\ldots,m\}$, from equation 20, we have
@@
Combining inequality equation 28, equation 29 and equation 30 yields
@@
Plugging inequalities equation 32 to equation 31, we get
@@
Taking the full expectation on both sides of equation 33 and adding the result of Lemma B.2, equation 14 and equation 13, we have
@@
Taking summation from $t=0$ to T1T-1 yields
@@
$\mathcal{C}=\max\{$ $8L_{f}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|,2\rho_{f}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|^{2},8\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|$
@@
By Jensen’s inequality, we can get $\mathbb{E}[\|\nabla\Phi_{1/\bar{\rho}}(\mathbf{{x}}_{\hat{t}})\|]\leq\mathcal{O}(\epsilon)$, where $\hat{t}$ is selected uniformly at random from $\{1,\ldots,T\}$. ∎
@@
Proof of Theorem D.3.
From equation 28, also denote $\bar{\mathbf{x}}_{t}:=\text{prox}_{\Phi/\bar{\rho}}(\mathbf{x}_{t})$ and consider change in the Moreau envelope, we have
@@
Adding above two estimation to equation 34, we have
@@
Adding above inequality to equation 35, we have
@@
Adding the result of Lemma B.2 and inequality equation 13, we get
@@
Taking summation from $t=0$ to T1T-1 yields
@@
$\mathcal{C}=\max\{$ $4L_{\nabla f}L_{g}\frac{1}{n}\sum_{i=1}^{n}\|u_{1,i}^{0}-g_{i}(\mathbf{x}_{0})\|^{2},8\frac{1}{m}\sum_{k=1}^{m}\|u_{2,k}^{0}-h_{k}(\mathbf{x}_{0})\|$